Comfort chair company manufacturers a standard recliner. During February, the firm's Assembly Department started production of 73,700 chairs. During the month, the firm completed 78,500 chairs, and transferred them to the Finishing Department. The firm ended the month with 11,200 chairs in ending inventory. There were 16,000 chairs in beginning inventory. All direct materials costs are added at the beginning of the production cycle and conversion costs are added uniformly throughout the production process. The FIFO method of process costing is used by Comfort. Beginning work in process was 20% complete as to conversion costs, while ending work in process was 80% complete as to conversion costs.

Beginning inventory:

      Direct materials                   $24,800

      Conversion costs                  $35,900

Manufacturing costs added during the accounting period:

      Direct materials                  $184,250

      Conversion costs                $278,058

1. What are the equivalent units of production for conversion costs IF the Weighted Average method had been used?

2. What are the direct materials equivalent units if FIFO is used?

3. What is the cost per unit for direct materials if FIO is used?

4. What is Total Cost per unit if FIFO is used?
Answer #1

1) Equivalent unit of conversion if WAC used = 78500+(11200*80%) = 87460 Units

2) Equivalent unit of material if FIFO used = 73700 Units

3) Cost per unit for direct material if FIFO used = 184250/73700 = 2.5

4) Equivalent unit of conversion = (16000*80%+62500+11200*80%) = 84260

Conversion Cost per equivalent unit = 278058/84260 = 3.3

Total cost per unit if FIFO used = 2.5+3.3 = 5.8